OpenStudy (anonymous):

A little confused about why the absolute value signs are needed after the following...

OpenStudy (anonymous):

\[\sqrt{a^2\cos^2x}\] \[=a|cosx|\]

OpenStudy (anonymous):

Because if it's being squared under the root, wouldn't either a negative or +ve value for either a or cosx keep the value under the root positive?

OpenStudy (trojanpoem):

yeah , It may have been -acosx then It was ^2 sqrt

OpenStudy (trojanpoem):

I think you confused between Cos(-x) and Abs

OpenStudy (trojanpoem):

so either acosx or -acosx